Tuesday, November 18: SWING Shift

From Hall A Wiki
Jump to: navigation, search

Go back to Daily Run Plan Daily Run Plans

  • General considerations
    • Vincent Sulkosky is the current Run Coordinator. Please contact me directly via RC phone 757-876-1787
    • Follow the generic Instructions for shift takers
    • Shift leaders: Please make sure to fill out the beam time accounting form and write a "Shift Summary" log entry at the end of your shift regardless if there is beam or not.
    • In general, technical staff should not be contacted by the shift crew - please call the RC first.
  • Instructions for this shift
    • The target operator is required to supervise the cryo-target.
    • Kinematics, DVCS Kin1: Ebeam = 7.3 GeV, PLHRS=1.1981 GeV, Theta_HRS=22.83, Theta_calo=12.35
    • Make sure to use the DVCS configuration for all runs.
    • Due to continued beam trips at 1C12 and the required work needed to fix the beam issues, Hall A will not receive CW the rest of today, though tune beam will be available for use.
    • Swing shift will begin in Controlled Access. Once work on the left HRS dipole controls is completed or Jack Segal is done for today, Hall A will go back to beam permit.
    • If Jack succeeds in ramping up the left Dipole to 2.71 GeV (about 937.3 A), then the rest of the spectrometer magnets should be set to 2.71 GeV. The regulation on the left Dipole should remain off. Remember to cycle only Q2 and Q3 to 1600 A first.